OpenStudy (anonymous):

integral of e to the power of -x^2..help to solve this plz...

hartnn (hartnn):

do you know about beta and gamma functions ?

hartnn (hartnn):

good try, but that is not correct. you cannot pull out variables out of integration, only constants. \(\huge \int\limits_{}^{}(\frac{ 1 }{ -2x })\frac{ -2x }{ 1 }e ^{-x ^{2}}\ne \frac{ 1 }{ -2x } \int\limits_{}^{}\frac{ -2x }{ 1 }e ^{-x ^{2}}\)

OpenStudy (anonymous):

thank you..but already solve it...by using taylor expansion..

OpenStudy (anonymous):

sry dont know what was i thinking.... Again Sorry for misleading you and thanks @hartnn